If you want to fit a linear regression line y=a+bx from the above data, then what is the value of 'a'?
1
Expert's answer
2021-07-28T06:50:00-0400
"\bar{x}=\dfrac{\sum_ix_i}{n}"

"\bar{y}=\dfrac{\sum_iy_i}{n}"

"SS_{xx}=\sum_i(x_i-\bar{x})^2=\sum_ix_i^2-n(\bar{x})^2"

"SS_{yy}=\sum_i(y_i-\bar{y})^2=\sum_iy_i^2-n(\bar{y})^2"

"SS_{xy}=\sum_i(x_i-\bar{x})(y_i-\bar{y})=\sum_ix_iy_i-n\bar{x}\bar{y}"

"b=\dfrac{SS_{xy}}{SS_{xx}}"

"a=\bar{y}-b\bar{x}"